# A Rapid Trajectory Optimization and Control Framework for Resource-Constrained Applications

### v1

@@ I Introduction
However, the additional degrees of freedom associated with the floating base of the robot manipulator renders the analysis of system dynamics and real-time control challenging [7]. Furthermore, if the contact and capture of the target satellite are not planned carefully, the resulting momentum transfer after the capture can cause severe stability issues [8]. Consequently, most of such missions have been considered to be at a very low Technology Readiness Level (TRL) due to limited onboard autonomy constrained by Size, Weight, and Power (SWaP) [9].
On the contrary, the recent advancements in consumer electronics and sensor technology have pushed the boundaries of autonomous in-space operations [10]. The entire process of qualification, verification, and validation of such novel hardware components and software packages have been greatly accelerated by the ease of availability of CubeSat manufacturers and launch providers [11]. The new form-factor of nano-satellites has also been aptly leveraged for earth observation [12], oceanography [13] and space weather monitoring [14] applications.
The Model-Predictive-Control (MPC) framework has been extensively studied for automatic control of industrial processes [15]. Although the framework has seen limited use in aerospace vehicles due to its stringent qualification, robustness and real-time requirements [16], there is an unequivocal pursuit to overcome some of these shortcomings about the complexity and computational cost. The advancements have helped bridge the gap between computationally intensive MPC and resource-constrained computers by exploiting the control structures to accelerate and compress solver algorithms [17]. This has also motivated recent studies to assess the computational requirements of a pragmatic MPC-based RPOD algorithm with the current state-of-the-art solvers [18].
Alongside the advances in available onboard hardware, many new computationally efficient algorithms have been developed to solve optimal control problems. In particular, pseudospectral methods are well-known for their accuracy and robustness [19]. Recent work using integral collocation, rather than the typical derivative methods, yields enhanced precision and significant runtime speedups [20, 21]. One of these methods, known as Integral Chebyshev Collocation (ICC), has been shown to be especially accurate, and does not require the use of an another (typically Legendre) interpolating polynomial [21].
To that end, this paper presents a pseudospectral-based MPC framework called MPC3: Model Predictive Control using Chebyshev Collocation. The main contribution of the paper, constrained quadratic programming optimization via orthogonal approximation using ICC, is introduced in Sec. II. Performance results for the proposed method are compared with existing solvers in Sec. III. Finally, Sec. IV demonstrates the effectiveness of the framework to achieve the desired objective with an application of multi-state guidance logic for a docking scenario of small satellites.
@@ II MPC3 Formulation
II-A Chebyshev Polynomial Approximation
Orthogonal polynomials play an important role in the fields of numerical analysis, approximation, and optimal control. A particularly attractive family of these orthogonal functions are the Chebyshev polynomials, which are defined over the domain $\tau\in[-1,1]$ as
$T_{n}(\tau)$ $=\cos\left(n\arccos\tau\right)$ (1)
Any function $f(t)$ can be mapped from the time domain $t\in[t_{0},t_{f}]$ to the computational domain $\tau\in[-1,1]$ over which the Chebyshev polynomials are defined with the linear transformation
$\tau$ $=\frac{2t-(t_{f}-t_{0})}{\Delta t},\quad\Delta t=t_{f}-t_{0}$ (2)
Thus, an approximation of the function at any point $f(\tau)$ can be constructed as a linear combination
$f(\tau)$ $=\sum_{i=0}^{n}T_{i}(\tau)\alpha=\bm{T}(\tau)\bm{\alpha}$ (3)
where $\bm{T}(\tau)=[T_{0}(\tau),T_{1}(\tau),\ldots,T_{n}(\tau)]\in\mathbb{R}^{1\times(n+1)}$ is a row vector of Chebyshev polynomials evaluated at $\tau$ and $\bm{\alpha}=[\alpha_{0},\alpha_{1},\ldots,\alpha_{n}]^{\top}\in\mathbb{R}^{(n+1)\times 1}$ is a column vector of unknown coefficients. Using a set of nodes $\bm{\tau}=[\tau_{0},\tau_{1},\ldots,\tau_{n}]^{\top}$, the approximation of the function $\bm{f}(\bm{\tau})=[f(\tau_{0}),f(\tau_{1}),\ldots,f(\tau_{n})]^{\top}$ is formed as
$\begin{bmatrix}f(\tau_{0})\\ f(\tau_{1})\\ \vdots\\ f(\tau_{n})\end{bmatrix}$ $=\begin{bmatrix}T_{0}(\tau_{0})&T_{1}(\tau_{0})&\cdots&T_{n}(\tau_{0})\\ T_{0}(\tau_{1})&T_{1}(\tau_{1})&\cdots&T_{n}(\tau_{1})\\ \vdots&\vdots&\ddots&\vdots\\ T_{0}(\tau_{n})&T_{1}(\tau_{n})&\cdots&T_{n}(\tau_{n})\end{bmatrix}\begin{bmatrix}\alpha_{0}\\ \alpha_{1}\\ \vdots\\ \alpha_{n}\end{bmatrix}$ (4)
$\bm{f}(\bm{\tau})$ $=\mathcal{T}(\bm{\tau})\bm{\alpha},\quad\mathcal{T}(\bm{\tau})\in\mathbb{R}^{(n+1)\times(n+1)}$ (5)
It has been shown that the optimal set of nodes $\bm{\tau}$ which minimize the interpolation error are the roots of the $(n+1)^{\text{th}}$ order Chebyshev polynomial [22], known as the Chebyshev-Gauss (CG) nodes.
II-B Integral Chebyshev Collocation
Chebyshev polynomials have been the subject of much work for solving initial value problems (IVPs) and boundary value problems (BVPs). Bai and Junkins developed a procedure known as Modified Chebyshev Picard Iteration (MCPI) to converge toward a solution for IVPs and BVPs [20, 23]. More recently, Peck proposed a concise procedure known as Integral Chebyshev Collocation (ICC) [21], the method of choice for this work. This approach is described hereafter for a generic second-order dynamic system.
The ICC procedure begins by transforming the differential equation(s) to the Chebyshev domain $\tau\in[-1,1]$. This is achieved simply via the chain rule by differentiating Eq. (2) with respect to $\tau$,
$\ddot{x}(t)$ $=\frac{\mathrm{d}^{2}x}{\mathrm{d}t^{2}}=\frac{\mathrm{d}^{2}x}{\mathrm{d}\tau^{2}}\frac{\mathop{}\!\mathrm{d}{\tau}^{2}}{\mathop{}\!\mathrm{d}{t}^{2}}=x^{\prime\prime}(\tau)\left(\frac{2}{\Delta t}\right)^{2}$ (7)
$x^{\prime\prime}(\tau)$ $=\left(\frac{\Delta t}{2}\right)^{2}\ddot{x}(t)$ (8)
The first derivative and state are then transformed as
$x^{\prime}(\tau)$ $=\left(\frac{\Delta t}{2}\right)\dot{x}(t)$ (9)
$x(\tau)$ $=x(t)$ (10)
Similarly, the initial conditions $\dot{x}_{0}$ and $x_{0}$ are cast to the computational domain as
$x^{\prime}(-1)$ $=\left(\frac{\Delta t}{2}\right)\dot{x}_{0},\quad x(-1)=x_{0}$ (11)
Now that the problem is represented in the Chebyshev domain, an $n^{\text{th}}$ order series of Chebyshev polynomials is used to represent the dynamics following Eq. (3).
@@ II-B Integral Chebyshev Collocation
The derivative representation is then formed by integrating this series from the initial time (in the Chebyshev domain) $\tau=-1$ to the current time $\tau$.
$\int_{-1}^{\tau}x^{\prime\prime}(\tau)\mathop{}\!\mathrm{d}{\tau}$ $=\int_{-1}^{\tau}\bm{T}(\tau)\bm{\alpha}\mathop{}\!\mathrm{d}{\tau}$ (13)
$x^{\prime}(\tau)-x^{\prime}(-1)$ $=\left[\int_{-1}^{\tau}\bm{T}(\tau)\mathop{}\!\mathrm{d}{\tau}\right]\bm{\alpha}$ (14)
The remaining bracketed vector integral can be taken term by term, where the integrals of the Chebyshev polynomials are given as [24, 25]
$\int T_{0}(\tau)\mathop{}\!\mathrm{d}{\tau}$ $=T_{1}(\tau)$ (15)
$\int T_{1}(\tau)\mathop{}\!\mathrm{d}{\tau}$ $=\frac{1}{4}\left(T_{2}(\tau)+T_{0}(\tau)\right)$ (16)
$\int T_{i}(\tau)\mathop{}\!\mathrm{d}{\tau}$ $=\frac{1}{2}\left(\frac{T_{i+1}(\tau)}{i+1}-\frac{T_{i-1}(\tau)}{i-1}\right),\quad i\geq 2$ (17)
The derivative is represented as
$x^{\prime}(\tau)$ $=\underbrace{\left[\bm{T}(\tau)\tilde{\mathcal{F}}-\mathbf{v}(-1)\right]}_{\equiv\bm{\beta}(\tau)}\bm{\alpha}+x^{\prime}(-1)$ (18)
where ~$\tilde{\mathcal{F}}\in\mathbb{R}^{(n+1)\times(n+1)}$ is a square matrix composed of integration weights on the super- and sub-diagonals, and $\mathbf{v}(-1)\in\mathbb{R}^{1\times(n+1)}$ is the evaluated lower integration bound [21].
Following the same procedure by integrating Eq. (18) yields the state approximation,
$x(\tau)$ $=\underbrace{\left[\bm{T}(\tau)\tilde{\mathcal{F}}^{2}-\mathbf{v}(-1)\tilde{\mathcal{F}}-\mathbf{v}(-1)(\tau+1)\right]}_{\equiv\bm{\gamma}(\tau)}\bm{\alpha}$
$\quad+x^{\prime}(-1)(\tau+1)+x(-1)$ (19)
Thus, the dynamics, derivative, and state are all approximated with a single set of polynomial coefficients $\bm{\alpha}$.
This procedure is then applied to CG nodes $\bm{\tau}$, leading to the following vector-matrix collocation formulation, which can be written succinctly using the $\bm{\beta}$ and $\bm{\gamma}$ shorthand defined above.
$\bm{x}^{\prime\prime}(\bm{\tau})$ $=\mathcal{T}(\bm{\tau})\bm{\alpha}$ (20)
$\bm{x}^{\prime}(\bm{\tau})$ $=\bm{\beta}(\bm{\tau})\bm{\alpha}+\bm{x}^{\prime}(-1)$ (21)
$\bm{x}(\bm{\tau})$ $=\bm{\gamma}(\bm{\tau})\bm{\alpha}+\bm{x}^{\prime}(-1)(\bm{\tau}+1)+\bm{x}(-1)$ (22)
Once an approximation order $n$ is chosen, $\mathcal{T}(\bm{\tau})$, $\bm{\beta}(\bm{\tau})$, and $\bm{\gamma}(\bm{\tau})$ are all constant, and the initial conditions are simply added to the matrix-vector product. This leads to a drastic improvement in computational efficiency since these matrices need to be computed once, which can be done a priori. To solve a linear system, a single matrix inversion is required, while nonlinear systems require an iterative method to converge on the coefficient values.
II-C $L_{2}$ Minimization
The objective of a typical MPC is to solve an optimization problem at each time instance for a specified control horizon and produce a control input that minimizes the objective function [26]. Unlike traditional discrete MPC, the cost function for the linear dynamics $\dot{\bm{x}}(t)=\bm{g}(\bm{t},\bm{x},\bm{u})$ is given as a continuous integral from the current time $t_{0}$ to some future time $t_{f}$. This is written as
$\min J$ $=\int_{t_{0}}^{t_{f}}\left(\bm{u}^{\top}\bm{W_{u}}\bm{u}+(\bm{x}-\bm{x}_{t})^{\top}\bm{W_{x}}(\bm{x}-\bm{x}_{t})\right)\mathop{}\!\mathrm{d}{t}$ (23)
where $\bm{u}\in\mathbb{R}^{m}$ is the control, $\bm{x},\bm{x}_{t}\in\mathbb{R}^{q}$ are the instantaneous and target state, and $\bm{W_{u}}\in\mathbb{R}^{m\times m}$ and $\bm{W_{x}}\in\mathbb{R}^{q\times q}$ are the diagonal weight matrices for the control and states, respectively. $\Delta t=t_{f}-t_{0}$ is now the time or prediction horizon associated with the problem.
For a generic double integrator $\bm{\ddot{x}}(t)=\bm{u}(t)$, the states and controls can both be represented by the unknown polynomial coefficients $\bm{\alpha}$. Without loss of generality, consider a single degree of freedom system, which can be written as
$x^{\prime\prime}(\tau)$ $=u(\tau)=\bm{T}(\tau)\bm{\alpha}$ (24)
@@ II-C $L_{2}$ Minimization
The problem is then converted to the Chebyshev domain as
$\min J$ $=\frac{\Delta t}{2}\int_{-1}^{1}\left[u(\tau)W_{u}u(\tau)\right.$
$\qquad\quad\qquad+(x(\tau)-x_{t})W_{x}(x(\tau)-x_{t})$
$\qquad\quad\qquad+\left.(x^{\prime}(\tau)-x_{t}^{\prime})W_{x^{\prime}}(x^{\prime}(\tau)-x_{t}^{\prime})\right]\mathop{}\!\mathrm{d}{\tau}$ (27)
Plugging in Eqs. (24)-(26) leads to a form that includes terms with quadratic, linear, and no dependence on the coefficients $\bm{\alpha}$. The terms independent of the decision variable are constant, and can be removed from the minimization problem [27]. However, this leaves a form which is not easily integrable, given the Chebyshev polynomial representations.
To evaluate this integral, a Gaussian quadrature rule is used, which approximates the integral as a linear combination of quadrature weights and the integrand evaluated at specific node points [28].
Using the CG nodes of Eq. (6), the weights are given as [29]
$w_{i}$ $=\frac{2}{n+1}\left[1-2\sum_{k=1}^{[(n+1)/2]}\frac{\cos(2k\theta_{i})}{4k^{2}-1}\right]$ (29)
$\theta_{i}$ $=\frac{2i-1}{2(n+1)}\pi,\quad\tau_{i}=\cos\theta_{i}$ (30)
Thus, the problem is finally written in the typical quadratic problem (QP) form
$\min J$ $=\frac{1}{2}\bm{\alpha}^{\top}H\bm{\alpha}+\bm{f}^{\top}\bm{\alpha}$ (31)
where the symmetric, positive definite matrix $H$ and vector $\bm{f}$ are given as
$\qquad\quad\left.+(\bm{\beta}(\tau_{i}))^{\top}W_{x^{\prime}}\bm{\beta}(\tau_{i})\right)\Bigg{]}$ (32)
$\qquad\quad\left.+W_{x^{\prime}}\bm{\beta}(\tau_{i})\left(x^{\prime}(-1)-x_{t}^{\prime}\right)\right)\Bigg{]}$ (33)
@@
Objective function $\sum_{i=0}^{p-1}\Big{(}\bm{u}^{T}_{k+i}\bm{W}_{u}\bm{u}_{k+i}\ +\left(\bm{y}_{k+i}-\bm{y}_{r}\right)^{T}\bm{W}_{y}\left(\bm{y}_{k+i}-\bm{y}_{r}\right)\Big{)}$ $\int_{t_{0}}^{t_{f}}\left(\bm{u}^{\top}\bm{W_{u}}\bm{u}+(\bm{x}-\bm{x}_{t})^{\top}\bm{W_{x}}(\bm{x}-\bm{x}_{t})\right)\mathop{}\!\mathrm{d}{t}$
Decision Variables $\bm{u_{k}}\in\mathbb{R}^{p\times m}$ $\bm{\alpha}\in\mathbb{R}^{(n+1)\times 1}$
# of Equality Constraints 00 $q$
@@
II-D State and Input Constraints

### v2

@@ I Introduction
However, the additional degrees of freedom (DoF) associated with the floating base of the robot manipulator renders the analysis of system dynamics and real-time control challenging [7]. Consequently, most of such missions have been considered to be at a very low Technology Readiness Level (TRL) due to limited onboard autonomy constrained by Size, Weight, and Power (SWaP) [8]. On the contrary, the recent advancements in consumer electronics and sensor technology have pushed the boundaries of autonomous in-space operations [9]. The process of qualification, verification, and validation of such novel hardware components and software packages have been greatly accelerated by the ease of availability of CubeSat manufacturers and launch providers [10].
The Model-Predictive-Control (MPC) framework has been extensively studied for automatic control of industrial processes [11]. Although the framework has seen limited use in aerospace vehicles due to its stringent qualification, robustness and real-time requirements [12]. There is an unequivocal pursuit to bridge the gap between computationally intensive MPC and resource-constrained computers by exploiting the control structures to accelerate and compress solver algorithms [13]. This has also motivated studies to assess the computational requirements of a pragmatic MPC-based RPOD algorithm [14].
Alongside the advances in available onboard hardware, many new computationally efficient algorithms have been developed to solve optimal control problems. In particular, pseudospectral methods are well-known for their accuracy and robustness [15]. Recent work using integral collocation, rather than the typical derivative methods, yields enhanced precision and significant runtime speedups [16, 17]. One of these methods, known as Integral Chebyshev Collocation (ICC) [17], has been shown to be especially accurate, and does not require the use of an another (typically Legendre) interpolating polynomial.
To that end, this paper presents a pseudospectral-based MPC framework called MPC3: Model Predictive Control via Chebyshev Collocation. The main contribution of the paper, constrained quadratic programming optimization via orthogonal approximation using ICC, is introduced in Sec. II. Performance results for the proposed method are compared with existing solvers in Sec. III. Finally, Sec. IV demonstrates the effectiveness of the framework to achieve the desired objective with an application of multi-state guidance logic for a docking scenario of small satellites.
@@ II MPC3 Formulation
II-A Integral Chebyshev Collocation
Orthogonal polynomials play an important role in the fields of numerical analysis, approximation, and optimal control. A particularly attractive family of these orthogonal functions are the Chebyshev polynomials,
$T_{n}(\tau)$ $=\cos\left(n\arccos\tau\right),\qquad\tau\in[-1,1]$ (1)
which have been the subject of much work for solving initial value problems (IVPs) and boundary value problems (BVPs). Bai and Junkins developed Modified Chebyshev Picard Iteration (MCPI) to converge toward a solution for IVPs and BVPs [16, 18]. More recently, Peck and Majji proposed a concise method called Integral Chebyshev Collocation (ICC) [17], the chosen framework for this paper, described briefly hereafter for a generic second-order system.
The ICC procedure begins by transforming the problem from the time domain $t\in[t_{0},t_{f}]$ to the computational domain $\tau\in[-1,1]$ over which the Chebyshev polynomials are defined. This is achieved with the linear transformation
$\tau$ $=\frac{2t-(t_{f}+t_{0})}{\Delta t},\qquad\Delta t=t_{f}-t_{0}$ (2)
The dynamics, derivative, state, and corresponding initial conditions are then cast to the computational domain as
$x^{\prime\prime}(\tau)$ $=\left(\frac{\Delta t}{2}\right)^{2}\ddot{x}(t)$ (3a)
$x^{\prime}(\tau)$ $=\left(\frac{\Delta t}{2}\right)\dot{x}(t),\quad x^{\prime}(-1)=\left(\frac{\Delta t}{2}\right)\dot{x}_{0}$ (3b)
$x(\tau)$ $=x(t),\quad x(-1)=x_{0}$ (3c)
Now that the problem is represented in the Chebyshev domain, an $n^{\text{th}}$ order series of Chebyshev polynomials is used to represent the dynamics,
@@ II-A Integral Chebyshev Collocation
where $\bm{T}(\tau)=[T_{0}(\tau),T_{1}(\tau),\ldots,T_{n}(\tau)]\in\mathbb{R}^{1\times(n+1)}$ is a row vector of Chebyshev polynomials evaluated at $\tau$ and $\bm{\alpha}=[\alpha_{0},\alpha_{1},\ldots,\alpha_{n}]^{\top}\in\mathbb{R}^{(n+1)\times 1}$ is a column vector of unknown coefficients. Integrating this expression of Chebyshev polynomials once or twice yields the derivative or state expression,
@@
where $\bm{\beta}(\tau),\bm{\gamma}(\tau)\in\mathbb{R}^{1\times(n+1)}$ are introduced as a shorthand to denote the first and second integration operators given in [17], respectively.
This procedure is then extended to the set of nodes $\bm{\tau}=[\tau_{0},\tau_{1},\ldots,\tau_{n}]^{\top}$ that minimize the interpolation error: the Chebyshev-Gauss (CG) nodes which the roots of the $(n+1)^{\text{th}}$ order Chebyshev polynomial [19],
Using the CG nodes $\bm{\tau}$ leads to a vector-matrix collocation formulation,
$\bm{x}^{\prime\prime}(\bm{\tau})$ $=\mathcal{T}(\bm{\tau})\bm{\alpha}$ (8a)
$\bm{x}^{\prime}(\bm{\tau})$ $=\bm{\beta}(\bm{\tau})\bm{\alpha}+\bm{x}^{\prime}(-1)$ (8b)
$\bm{x}(\bm{\tau})$ $=\bm{\gamma}(\bm{\tau})\bm{\alpha}+\bm{x}^{\prime}(-1)(\bm{\tau}+1)+\bm{x}(-1)$ (8c)
where $\mathcal{T}(\bm{\tau}),\bm{\beta}(\bm{\tau}),\bm{\gamma}(\bm{\tau})\in\mathbb{R}^{(n+1)\times(n+1)}$ are matrix versions of the previously defined vectors. Once the approximation order $n$ is chosen, these matrices are constant and can be computed a priori, leading to a drastic improvement in computational efficiency. To solve a linear system, the coefficients can be found via a single matrix inversion, while nonlinear systems require an iterative method.
@@
Objective function $\sum_{i=0}^{p-1}\Big{(}\bm{u}^{T}_{k+i}\bm{W}_{u}\bm{u}_{k+i}\ +\left(\bm{y}_{k+i}-\bm{y}_{r}\right)^{T}\bm{W}_{y}\left(\bm{y}_{k+i}-\bm{y}_{r}\right)\Big{)}$ $\int_{t_{0}}^{t_{f}}\left[\bm{u}^{\top}\bm{W_{u}}\bm{u}+(\bm{x}-\bm{x}_{t})^{\top}\bm{W_{x}}(\bm{x}-\bm{x}_{t})\right]\mathop{}\!\mathrm{d}{t}$
Decision Variables $\bm{u_{k}}\in\mathbb{R}^{p\times m}$ $\chi=\left[\bm{\alpha},\varepsilon\right]^{\top}\in\mathbb{R}^{(q(n+1)+1)\times 1}$
# of Equality Constraints 00 $2(n+2)\times q$
@@
II-B $L_{2}$ Minimization
The objective of a typical MPC is to solve an optimization problem at each time instance for a specified control horizon and produce a control input that minimizes the objective function [20]. Unlike traditional discrete MPC, the cost functional for the linear dynamics $\dot{\bm{x}}(t)=\bm{g}(\bm{t},\bm{x},\bm{u})$ is given as a continuous integral from the current time $t_{0}$ to some future time $t_{f}$,
$J$ $=\int_{t_{0}}^{t_{f}}\left[\bm{u}^{\top}\bm{W_{u}}\bm{u}+(\bm{x}-\bm{x}_{t})^{\top}\bm{W_{x}}(\bm{x}-\bm{x}_{t})\right]\mathop{}\!\mathrm{d}{t}$ (9)
where $\bm{u}\in\mathbb{R}^{m}$ is the control, $\bm{x},\bm{x}_{t}\in\mathbb{R}^{q}$ are the instantaneous and target state, and $\bm{W_{u}}\in\mathbb{R}^{m\times m}$ and $\bm{W_{x}}\in\mathbb{R}^{q\times q}$ are the diagonal weight matrices for the control and states, respectively. The time or prediction horizon associated with the problem is now $\Delta t$ from Eq. (2).
Hereafter, we consider problems which can be written or linearized in the form of a double integrator, $\bm{\ddot{x}}(t)=\bm{u}(t)$, the benefit of which is that both the states and controls can be represented with the same set of unknown polynomial coefficients. For the sake of brevity, a 1-DoF double integrator will be used in this derivation, but the procedure easily extends to multi-DoF problems by using another set of coefficients for each additional state, as is done in Sec. IV. The cost functional is then converted to the Chebyshev domain as
$J$ $=\frac{\Delta t}{2}\int_{-1}^{1}\big{[}u(\tau)W_{u}u(\tau)+(x(\tau)-x_{t})W_{x}(x(\tau)-x_{t})$
$\qquad\quad\qquad+(x^{\prime}(\tau)-x_{t}^{\prime})W_{x^{\prime}}(x^{\prime}(\tau)-x_{t}^{\prime})\big{]}\mathop{}\!\mathrm{d}{\tau}$ (10)
where the integrand includes terms with quadratic, linear, and no dependence on the coefficients $\bm{\alpha}$. The terms independent of the decision variable are constant, and can be removed from the minimization problem [21], but the remaining integrand expression is not well-posed given the Chebyshev polynomial representations of the state and derivative.
To evaluate this integral, a Gaussian quadrature rule is applied to approximate the integral,
where $w_{i}$ are the quadrature weights [22] and $g(\tau_{i})$ are the integrand evaluated at the CG nodes. Thus, the problem is written in the typical quadratic problem (QP) form
$\min J$ $=\frac{1}{2}\bm{\alpha}^{\top}H\bm{\alpha}+\bm{f}^{\top}\bm{\alpha}$ (12)
where the matrix $H\succ 0$ and vector $\bm{f}$ are given as
$\qquad\quad\left.+(\bm{\beta}(\tau_{i}))^{\top}W_{x^{\prime}}\bm{\beta}(\tau_{i})\right)\Bigg{]}\in\mathbb{R}^{(n+1)\times(n+1)}$ (13a)
$\qquad\quad\left.+W_{x^{\prime}}\bm{\beta}(\tau_{i})\left(x^{\prime}(-1)-x_{t}^{\prime}\right)\right)\Bigg{]}\in\mathbb{R}^{(n+1)\times 1}$ (13b)
II-C State and Input Constraints